- PSL had stayed on overnight. There was an EQ (M 4.6 near Costa Rica) which showed up on the Seis BLRMS, and I noticed that several optics were reporting Oplev spots off their QPDs (I had just centered these yesterday). So I did a quick alignment check:
- IMC was readily locked
- After moving test mass bias sliders to bring Oplev spots back to the center, the EX and EY green beams were readily locked to a TEM00 mode
- IR flashes could be seen in TRX and TRY (though their levels are low, since we are operating with 1/10th the nominal power
- The IP-POS QPD channels were reporting a "segmentation fault" so I keyed the c1iscaux crate and they came back. Still the QPD was reporting a low SUM value, but this too is because of the lower power. Conveniently, there was an ND2.0 filter in the beam path on a flip mount which I just flipped out of the way for the low-power tracking.
- Then, PSL and green shutters were closed and Oplev loops were disengaged.
- Checked that we have an RGA scan from today
- During the walkthrough to check the jam nuts, Chub noticed that the outer nuts on the bellows between the OMC chamber and the IMC chamber were loose to the finger! He is tightening them now and checking the remaining jam nuts. AFAIK, Steve made it sound like this was always a formality. Should we be concerned? The other jam nuts are fine according to Chub.
- We valved off the pumpspool from the main volume and annuli, and started letting Nitrogen into the main volume at ~1045am.
- Started letting instrument grade air into the main volume at ~1130am. We are aiming for a pressure increase of 3 torr/min
- 4 cylinders of dry air were exhausted by ~330pm. It actually looks like we over-pressured the main volume by ~20torr - this is bad, we should've stopped the air inletting at 700 psi and then let it equilibriate to lab air pressure.
- At some point during the vent, the main volume pressure exceeded the working range of the cold cathode gauge CC1. It reports "Current Fail" on its LED display, which I'm assuming meant it auto-shutoff its HV to protect itself, Jon tells me the vacuum code isn't responsible for initiating any manual shutoff.
- A new vacuum state was added to reflect these conditions (pumpspool under vacuum, main volume at atmosphere).
- The annuli remain under vacuum for now. Tomorrow, when we remove the EY door, we will vent the EY annulus.
IMC was locked, MC2T ~ 1200cts after some alginment touch ups. The test mass oplevs indicate some drift, ~100urad. I didn't realign them.
The EY door removal will only be done tomorrow. I will take some free-swinging ETMY data today (suspension was kicked at 1241919438) to see if anything has changed (it shouldn't have). I need to think up a systematic debugging plan in the meantime. |